OpenStudy (anonymous):

solve each system by substitution 4x-3y=12 x-2y=3

OpenStudy (anonymous):

x = 3 + 2y 4(3 + 2y) - 3y = 12 12 + 8y - 3y = 12 12 + 5y = 12 5y = 0 y = 0 x - 2(0) = 3 x = 3 Hopefully I didn't mess anything up.